OpenStudy (anonymous):

If 3x=2(5-2x),then x=?

OpenStudy (anonymous):

x=10/7

OpenStudy (uri):

Solve the paretheses first 3x=10-4x Bring 4x left side 3x+4x=10 7x=10 divide 10 with 7 x=10/7

OpenStudy (anonymous):

3x=2(5-2x) i.e. 3x= 10 - 4x i.e. 3x+4x=10 i.e. 7x= 10 i.e. x=10/7